CONCENTRATION, MATRIX, AND WAVELENGTH DEPENDENCE IN THE PHOTOLYSIS EFFICIENCY OF MATRIX-ISOLATED BIACETYL

C. A. BAUMANN, Department of Chemistry, The University of Scranton, Scranton, PA 18510-4626.

The ultraviolet irradiation of matrix-isolated biacetyl (C_{4}H_{6}O_{2}) results in the formation of a complex of transmethylhydroxycarbene (CH_{3}COH) and ketene (CH_{2}CO). Studies in this laboratory using pulsed sources have confirmed that these products arise from a two-photon process: the first photon (λ < 450 nm) places the molecule in the Ã ^{1}A_{u} state, followed by relaxation to the long-lived ã ^{3}A_{u} state. Triplet-triplet excitation(λ < 550 nm) from this state leads to the observed products. The apparent quantum efficiency for the formation of these products was measured as a function of matrix (N_{2}, O_{2}, Ar, Kr, and Xe), dilution ratio for the parent compound, and wavelength of irradiation (using a continuous source). For all matrices, the apparent efficiency decreases with increasing C_{4}H_{6}O_{2} dilution, indicating that the products may arise from a bimolecular process. No such effect was seen in C_{4}D_{6}O_{2} experiments, which exhibit much lower levels of efficiency at all dilutions. The matrix has an effect on the high-dilution efficiency (N_{2}, O_{2} > Kr, Xe > Ar), which suggests that the matrix participates in the H-transfer process. Increasing the wavelength of irradiation from 436 to 450 nm results in an increase in the measured efficiency. At 436 nm, the reverse intersystem crossing (ã ^{3}A_{u} → Ã ^{1}A_{u}) channel is available to drain the triplet population. At 450 nm, this channel is no longer available, and the resulting increase in the steady-state triplet population leads to the observed enhancement.